# Language Models Do Not Follow Occam’s Razor: A Benchmark for Inductive and Abductive Reasoning

### v1

@@ Abstract
Reasoning is a core capability in artificial intelligence systems, for which large language models (LLMs) have recently shown remarkable progress. However, most work focuses exclusively on deductive reasoning, which is problematic since other types of reasoning are also essential in solving real-world problems, and they are less explored. This work focuses on evaluating LLMs’ inductive and abductive reasoning capabilities. We introduce a programmable and synthetic dataset, InAbHyD (pronounced in-a-bid), where each reasoning example consists of an incomplete world model and a set of observations. The task for the intelligent agent is to produce hypotheses to explain observations under the incomplete world model to solve each reasoning example. We propose a new metric to evaluate the quality of hypotheses based on Occam’s Razor. We evaluate and analyze some state-of-the-art LLMs. Our analysis shows that LLMs can perform inductive and abductive reasoning in simple scenarios, but struggle with complex world models and producing high-quality hypotheses, even with popular reasoning-enhancing techniques such as in-context learning and RLVR.  [^1] All data and generation code are freely available at https://github.com/byrantwithyou/inabhyd/.
@@ 1 Introduction
Reasoning is a hallmark of human intelligence. Peirce categorizes reasoning into deductive, inductive, and abductive reasoning  (Burch, 2024). Deductive reasoning draws a specific conclusion from a set of premises. For example, given “All humans are moral” and “Socrates is a human”, we can conclude that “Socrates is moral”. Inductive reasoning aims to identify a general principle from multiple individual cases. For example, if we see many cute cats, we may inductively reason that all cats are cute. Abductive reasoning seeks to produce hypotheses to explain observations. For example, if we hear a bird sound outside, we may abductively reason that there is a bird. The goal in both types of reasoning is to find explanations to explain observations (inductive reasoning focuses on inducing a more general rule, while abductive reasoning focuses on finding a ground fact or axiom).
The composition of inductive and abductive reasoning is ubiquitous in real-world scenarios. For example, given (i) Amy has a skin rash and leg pain and has been diagnosed with a new unknown virus. (ii) Sam has a skin rash and leg pain and has been diagnosed with the same new unknown virus. (iii) Tom also has a skin rash and leg pain. We can first induce the rule: The new unknown virus causes skin rashes and leg pain, and then we can abduce that Tom is infected with this new virus.
A fundamental difference between deductive and inductive or abductive reasoning is that deductive reasoning is deterministic. The conclusion in deductive reasoning is deterministically correct if the proof is sound and the premises are correct. However, inductive and abductive reasoning are inherently probabilistic. Technically, there may exist many conclusions, some more probable (high-quality) and some less probable (low-quality). For example, in the above bird sound case, a less probable conclusion is that someone outside is playing the bird sound on a speaker. In this regard, we view inductive and abductive reasoning as the inverse processes of deductive reasoning and study them through a unified lens. In many real-world scenarios, all three reasoning modalities are indispensable. Deductive reasoning is widely used in mathematical problem-solving, inductive reasoning is commonly used in scientific research and discovery, and abductive reasoning is often employed in medical diagnosis. Knowing and applying the principles of parsimony / Occam’s Razor is essential for LLMs to be utilized in areas such as scientific discovery, which require inductive and abductive reasoning, yet this is often overlooked. Physicists, for example, strive for the simplest explanations to account for as many observations as possible. According to our findings, LLMs struggle to generate high-quality hypotheses that match human-level intelligence. To the best of our knowledge, prior studies of inductive and abductive reasoning did not consider parsimony / Occam’s Razor in their evaluation.
The high-level research question we want to answer in this paper is: Do LLMs reason inductively and abductively? If so, how well? There is a significant body of work on evaluating the deductive reasoning capabilities of LLMs  (Saparov and He, 2023; Saparov et al., 2023; Liu et al., 2020; Tafjord et al., 2021; Han et al., 2024), as well as techniques to improve deductive reasoning capabilities, such as in-context learning and reinforcement learning with verifiable reward, RLVR  (Lambert et al., 2024; Su et al., 2025). However, there is relatively little exploration into the inductive and abductive reasoning capabilities of large language models (LLMs). This work systematically evaluates whether and to what extent LLMs can reason abductively and inductively. We also assess whether popular reasoning-enhancing techniques, such as in-context learning  (Zhou et al., 2022; Dong et al., 2024) and RLVR  (Lambert et al., 2024; Su et al., 2025), which primarily aim to improve deductive reasoning capabilities, can also benefit inductive and abductive reasoning.
@@
Figure 1: An example of a reasoning question in InAbHyD . The upper left side is the world model. Note that the dashed arrow means the subtype relationship All rats are mammals is hidden and needs to be inferred (i.e., a hypothesis to propose). Each node is a concept (e.g., cat), with its properties and members. Each edge denotes a subtype relationship (e.g., All cats are carnivores). On the upper right-hand side are observations in this reasoning example and hypotheses (ground truth answers for this reasoning example). Each hypothesis explains observations of the same color (e.g., the hypothesis All mammals are hairy explains the observations Sam is hairy; Alice is hairy; Bob is hairy). On the lower side is the natural language description of this reasoning example. Note that this example uses a real world model for illustrative purposes. In our dataset, we use fictional world models. A more detailed explanation of this example is given in Appendix A.
@@
We create a programmable and synthetic dataset, InAbHyD (Inductive and Abductive Hypothesis Discovery), to test LLMs’ inductive and abductive reasoning capabilities. Each reasoning example has a fictional incomplete world model in the form of an ontology tree and a set of observations, represented in first-order logic. To answer each reasoning example, LLMs must produce hypotheses to explain observations under the incomplete world model. We generate synthetic examples of fictional world models to mitigate training data contamination. The ontology tree’s height characterizes the difficulty of the world model and the corresponding reasoning example. [^2] See Figure 1 for an example and Appendix A for more explanations for Figure 1 and more examples in Appendix B.
@@
We evaluate our dataset on five widely-used LLMs: Llama3-70B  (Dubey et al., 2024) , Gemma3-27B  (Kamath et al., 2025), DeepSeek-V3  (DeepSeek-AI et al., 2024), GPT-4o  (OpenAI, 2023) , and DeepSeek-R1-Distill-Llama-70B  (DeepSeek-AI et al., 2025). For each LLM, we test whether it can produce valid and high-quality hypotheses. We also test whether popular reasoning-enhancing techniques, such as in-context learning and RLVR, which primarily were meant to improve deductive reasoning, can also benefit inductive and abductive reasoning. Our findings can be summarized as follows:
@@
    The accuracy drops significantly when the ontology tree structure is complex and there are multiple hypotheses.
@@
    In some cases, LLMs may produce valid hypotheses (hypotheses which can explain all observations); but fail to produce high-quality hypotheses.
@@
    In general, we observe that reasoning-enhancing techniques for deductive reasoning only moderately benefits inductive and abductive reasoning.
@@ 2 Related Work
There is no consensus on whether LLMs can perform general deductive reasoning or are stochastic parrots (Bender et al., 2021). Researchers have designed different datasets and evaluation methodologies to answer this question  (Saparov et al., 2023; Saparov and He, 2023; Liu et al., 2020; Tafjord et al., 2021; Han et al., 2024). LogiQA  (Liu et al., 2020) is a multiple-choice QA dataset to test multiple types of deductive reasoning of language models. However, it only attends to label accuracy (the correctness of the final answer) without considering the correctness of intermediate proof steps. In contrast, PrOntoQA  (Saparov and He, 2023) and PrOntoQA-OOD  (Saparov et al., 2023) evaluate both label accuracy and step-by-step proof correctness. They can synthetically generate proofs with different heights, widths, and deduction rules. Our work is similar to PrOntoQA and ProntoQA-OOD in spirit. We both use a first-order ontology tree as the world model. However, PrOntoQA and PrOntoQA-OOD do not evaluate inductive and abductive reasoning.
Evaluating non-deductive reasoning capabilities.
[…]
Dataset Controls for complexity of world models Natural language description Inductive reasoning Abductive reasoning Evaluation of hypotheses quality Multiple hypotheses
ART (Zhao et al., 2023) non-quantitative
[…]
AbductionRules (Young et al., 2022)
Wino-Why (Zhang et al., 2020)
[…]
UniADILR (Sheng et al., 2025)
Mirage (Li et al., 2025)
Case2Code (Shao et al., 2024)
InAbHyD (ours)
Table 1 summarizes the similarities and differences between our work and others on evaluating non-deductive reasoning capabilities in both the pre- and LLM eras  (Young et al., 2022; Hua et al., 2025; Zhao et al., 2023; Yang et al., 2024; Zhang et al., 2021). Our work differs from others in the following ways. First, we focus on inductive and abductive reasoning and their interactions with deductive reasoning, while other works focus only on a single reasoning modality. Second, our reasoning problem is built upon a fictional world model with natural language descriptions. The fictional world model avoids training data contamination compared with ART  (Zhao et al., 2023), where the problem is directly taken from a web corpus; the natural language descriptions of the reasoning problem render the evaluation more practical compared with symbolic reasoning, such as in InductiveBench  (Hua et al., 2025). Third, our dataset is programmable with highly-controllable difficulty parameters, allowing us to evaluate LLMs’ reasoning capabilities under various scenarios.
@@
A large body of work has focused on improving the (deductive) reasoning capabilities of LLMs and related downstream tasks such as mathematics and coding  (Li, 2025; Wei et al., 2022; Yao et al., 2023; Wang et al., 2023; Su et al., 2025; Lambert et al., 2024; Zhou et al., 2022; Dong et al., 2024). The first line of work aims to elicit the emergent capabilities of LLMs by different formats of prompts, such as chain-of-thought (CoT)  (Wei et al., 2022), tree-of-thought (ToT)  (Yao et al., 2023), and self-consistency  (Wang et al., 2023). The second line of work leverages in-context learning by providing few-shot examples to LLMs. The third line of work views LLM token generation as a reinforcement learning problem and utilizes verifiable reward functions (reinforcement learning with verifiable reward, RLVRLambert et al. 2024; Su et al. 2025) to steer LLMs to improved reasoning. The fourth line of work scales LLMs’ test time to increase the accuracy of reasoning problems. It is worth noting that most techniques discussed above focus mainly on deductive reasoning. This work investigates whether the above methods benefit inductive and abductive reasoning.
@@ 3 Approach
In this section, we describe how we generate the dataset and evaluate LLMs on this dataset. At a high level, the dataset consists of synthetically-generated reasoning examples in natural language. The question description is a user prompt, and the format specification is a system prompt to LLMs. We parse LLMs’ output to evaluate their inductive and abductive reasoning capabilities on metrics such as accuracy. We acknowledge that the potential space of inductive and abductive reasoning examples is vast, so we restrict ourselves to what is expressible in first-order logic. By composing the inductive and abductive reasoning steps in Table 2, our evaluation framework can readily cover all inductive and abductive reasoning questions in first-order logic.
@@
    Second, we traverse the ontology tree and randomly mark some axioms as hidden. Each hidden axiom is the ground truth for hypotheses.
@@
    Third, for each hidden axiom, we generate observations so that LLMs have enough information to produce ground truth hypotheses.
@@
    Fourth, we translate the incomplete world model and observations into natural language.
Subtask Reasoning type tested World model Observations Ground truth hypothesis Proof trees for observations
@@
Table 2: An overview of various subtasks in InAbHyD and their details in the simplest case (a height-1 ontology tree with a single hypothesis). Here o denotes an ontology concept (e.g, cat) and o(A) denotes the membership (e.g., cat(Sally) means Sally is a cat). p denotes a property (e.g, cute) and p(A) denotes A has property p (e.g., cute(Sally) means Sally is cute). We only show the proof tree for one observation for simplicity; other observations can be derived similarly. We show hypotheses green and axioms in world models purple in the proof tree.
@@
    Strong accuracy: Hypotheses are strongly correct if they exactly match ground truth hypotheses.
@@
    Hypothesis quality: A quantitative measurement of the quality of the LLMs’ produced hypotheses based on the principles of parsimony.
@@
World model generation
@@
In our approach, the world model is an ontology tree in first-order logic. Each node in the ontology tree is a concept (e.g. cat) with properties (e.g. All cats are cute, or equivalently in logical form $\forall x(\texttt{cat}(x)\rightarrow\texttt{cute}(x))$) and members (e.g. Amy is a cat, or equivalently in logical forms, cat(Amy)). Each edge in the ontology tree denotes a subtype relationship (e.g., the ragdoll node being a child of the cat node implies Each ragdoll is a cat, or equivalently in logical form, $\forall x(\texttt{ragdoll}(x)\rightarrow\texttt{cat}(x))$). We first generate the topology of the ontology tree with the height as a controllable parameter. Each non-leaf node has two or three child nodes. The height of the ontology tree is a variable that characterizes the complexity of the world model. We then make two passes of the tree in level order traversal. In the first pass, for each node, we allocate a concept name, three properties, [^4] The properties can also be negated properties, e.g., All cats are not slow. and three members. After the first pass, we have a complete world model with three types of axioms: (1) properties (e.g., All cats are cute), (2) membership relations (e.g., Amy is a cat), (3) subtype relations (e.g., Each ragdoll is a cat). In the second pass, we randomly mark some axioms as hidden. These axioms are the ground truth hypotheses. After the second pass, we have an incomplete world model.
Observation generation

### v2

# Do Language Models Follow Occam’s Razor? An Evaluation of Parsimony in Inductive and Abductive Reasoning

@@ Abstract
Non-deductive reasoning, encompassing inductive and abductive reasoning, is essential in addressing complex real-world questions. One key feature of inductive and abductive reasoning is that there are many valid hypotheses; the simplest ones (those that adhere to Occam’s Razor) are often most useful. However, this aspect is ignored in recent work that evaluates the non-deductive reasoning capabilities of large language models (LLMs). This work fills this gap, focusing on understanding whether the inductive and abductive reasoning capabilities of LLMs adhere to Occam’s Razor, while also examining the correctness of their reasoning. To accomplish this goal, we introduce a framework to synthetically generate reasoning questions that (a) require inductive reasoning and abductive reasoning simultaneously; (b) is readily extended to produce any abductive/inductive reasoning question expressible in first-order logic. The task for the intelligent agent is to produce hypotheses to explain observations under a given world model. We also propose a new automated metric to assess whether hypotheses quantitatively adhere to Occam’s Razor; those hypotheses that are correct and simplest are considered high-quality. Our findings on state-of-the-art LLMs suggest that LLMs can perform inductive and abductive reasoning in simple scenarios, but struggle with complex world models and with producing high-quality hypotheses, even with popular reasoning-enhancing techniques such as in-context learning and RLVR.
[…]
Do Language Models Follow Occam’s Razor?
An Evaluation of Parsimony in Inductive and Abductive Reasoning
Yunxin Sun  and Abulhair Saparov {sun1114, asaparov}@purdue.edu Purdue University
@@ 1 Introduction
Reasoning is a hallmark of intelligence. Peirce categorizes reasoning into deductive, inductive, and abductive reasoning  (peirce1878deduction; cat), all of which are indispensable in real-world scenarios. Deductive reasoning draws a conclusion from a set of premises. Inductive reasoning aims to identify a general principle from multiple individual cases. Abductive reasoning seeks to produce hypotheses to explain observations. See Appendix A for examples of different types of reasoning and for further background.
@@
Figure 1: An example of a question requiring inductive and abductive reasoning. Note that the dashed arrow means the subtype relationship All rats are mammals is hidden and needs to be inferred (i.e., a hypothesis to propose). This example uses a real-world model for illustrative purposes. In our dataset, we use fictional world models.
There is a significant body of work on evaluating the deductive reasoning capabilities of LLMs  (prontoqa; prontoood; logiqa; proofwriter; follo), as well as techniques to improve deductive reasoning capabilities, such as in-context learning and reinforcement learning with verifiable rewards, RLVR  (tulu; RLVRmultiple). However, there is relatively little exploration into the inductive and abductive reasoning capabilities of large language models (LLMs). A fundamental difference between deductive and inductive or abductive reasoning is that deductive reasoning is deterministic. However, inductive and abductive reasoning are inherently probabilistic. There may exist many hypotheses, some more probable and some less probable. For example, if we hear a bird sound, a more probable hypothesis is that there is a bird outside; while a less probable hypothesis is that someone outside is playing the bird sound on a speaker. Among these possible hypotheses, the simplest ones are often more useful, and we view them as more probable. This principle is known as Occam’s Razor. We call hypotheses that adhere to Occam’s Razor high-quality hypotheses. Knowing and applying Occam’s Razor is essential in areas such as scientific discovery, which require inductive and abductive reasoning, yet this is often overlooked. Physicists, for example, strive for the simplest explanations to account for as many observations as possible.
The high-level research question we aim to answer is: Do LLMs reason inductively and abductively? If so, how well do they adhere to Occam’s Razor? This work systematically evaluates whether and to what extent LLMs can reason inductively and abductively and adhere to Occam’s Razor. We also assess whether popular reasoning-enhancing techniques, such as in-context learning  (ICL1; ICL2) and RLVR  (tulu; RLVRmultiple), which primarily aim to improve deductive reasoning capabilities, can also improve inductive and abductive reasoning.
@@
To answer this research question, we introduce a framework to synthetically generate reasoning questions that (a) contain inductive reasoning and abductive reasoning simultaneously; (b) is readily extended to produce any inductive/abductive reasoning question expressible in first-order logic. Each reasoning example has a fictional incomplete world model in the form of an ontology tree and a set of observations, represented in first-order logic. To answer each reasoning example, LLMs must produce hypotheses to explain observations under the incomplete world model. We generate synthetic examples of fictional world models to mitigate the contamination of training data. The ontology tree’s height characterizes the difficulty of the world model and the corresponding reasoning example. [^1] Refer to Figure 1 and Appendix C for examples. To assess whether LLMs adhere to Occam’s Razor quantitatively, we also introduce an automated metric to quantify the quality of each hypothesis (i.e., how well each hypothesis adheres to Occam’s Razor). We also introduce a new benchmark dataset, InAbHyD(Inductive and Abductive Hypothesis Discovery), generated under this framework. This dataset contains over 2K reasoning questions with different difficulty parameters. The dataset also includes the generation code, which can, in principle, be used to generate an unbounded number of reasoning questions.
@@
We evaluate our dataset on seven widely-used LLMs: GPT-5.4 (gpt5), o3 (o3), Llama3-70B  (llama) , Gemma3-27B  (gemma), DeepSeek-V3  (v3), GPT-4o  (gpt4) , and DeepSeek-R1-Distill-Llama-70B  (r1). For each LLM, we test whether it can produce valid and high-quality hypotheses. We also test whether popular reasoning-enhancing techniques, such as in-context learning and RLVR, which primarily were meant to improve deductive reasoning, can also improve inductive and abductive reasoning. Our findings can be summarized as follows:
@@
    The accuracy drops significantly when the ontology tree structure is complex and contains multiple hypotheses.
@@
    In most cases, LLMs may produce valid hypotheses (which can explain all observations); but fail to produce high-quality hypotheses (which adhere to Occam’s Razor).
@@
    In general, we observe that reasoning-enhancing techniques moderately improves inductive and abductive reasoning.
@@ 2 Related Work
Evaluating non-deductive reasoning capabilities.
Dataset Controls for complexity of world models Natural language description Tests for inductive reasoning Test for abductive reasoning Evaluation of hypotheses quality Multiple hypotheses
ART (anll) non-quantitative
[…]
AbductionRules (abductionrule)
Wino-Why (winowhy)
[…]
UniADILR (Uni)
LogiEval (logieval)
MME-Reasoning (mme_reasoning)
InAbHyD(ours)
Table 1: Comparison of existing datasets to evaluate inductive and abductive reasoning capabilities.
Table 1 summarizes the similarities and differences between our work and others on evaluating non-deductive reasoning capabilities of both LLMs and earlier models. Representative works on abductive reasoning evaluation, such as ART (anll), AbductionRules (abductionrule), and Wino-Why (winowhy), craft abductive reasoning questions from web sources, and ask LMs to explain everyday events. However, this approach confounds the problem of commonsense knowledge acquisition with the problem of abductive reasoning. In contrast, our framework provides all required facts in the context of each prompt, thereby evaluating LM’s abductive reasoning capabilities in a more controlled fashion. Representative works on inductive reasoning evaluation, such as InductiveBench (inductivebench) and SyGus (sygus) focus exclusively on highly-structured domains such as string translation and program synthesis. However, natural language-based inductive reasoning is also essential and is underexplored.
@@
There is no consensus on whether LLMs can perform general deductive reasoning or are stochastic parrots (parrot). Researchers have designed different datasets and evaluation methodologies to answer this question  (prontoood; prontoqa; logiqa; proofwriter; follo). LogiQA  (logiqa) is a multiple-choice QA dataset to test multiple types of deductive reasoning of language models. However, it focuses solely on label accuracy (the correctness of the final answer) without considering the correctness of intermediate proof steps. In contrast, PrOntoQA  (prontoqa; prontoood) evaluates both label accuracy and step-by-step proof correctness. They can synthetically generate proofs with different heights, widths, and deduction rules. Our work is similar to PrOntoQA in spirit. We both use a first-order ontology tree as the world model. However, PrOntoQA does not evaluate non-deductive reasoning.
@@
A large body of work has focused on improving the (deductive) reasoning capabilities of LLMs and related downstream tasks such as mathematics and coding  (ttt1; CoT; ToT; consistency; RLVRmultiple; tulu; ICL1; ICL2). The first line of work aims to elicit the emergent capabilities of LLMs using different prompt formats, such as chain-of-thought  (CoT; CoT), tree-of-thought  (ToT; ToT), self-consistency  (consistency), and in-context learning. The second line of work post-trains LLMs with reinforcement learning with verifiable reward (RLVR)  tulu; RLVRmultiple to improve LLM reasoning. Since non-deductive reasoning is inherently ambiguous, it is difficult to define a verifiable reward in non-deductive settings. The third line of work uses neuro-symbolic approaches to enhance the efficiency and reliability of LLM reasoning (ns1; ns2; ns3). However, they are either not directly applicable to unstructured natural language input or designed for deductive reasoning applications. This work investigates whether the above methods benefit inductive and abductive reasoning.
@@ 3 Approach
3.1 Overview
In this section, we describe how we generate the dataset and evaluate LLMs. The dataset consists of synthetically-generated reasoning examples in natural language. Figure 2 illustrates the pipeline for generating each question. We evaluate LLM’s inductive and abductive reasoning capabilities on metrics such as accuracy and their adherence to Occam’s Razor. By composing inductive and abductive reasoning steps shown in Table 2, our evaluation framework can readily be extended to produce any abductive/inductive reasoning question expressible in first-order logic.
Step 1 (§3.2) Generate the world model Step 2 (§3.3) Generate observations Step 3: Naturalistic Question Generation (§3.4) Convert theories and observations into natural language using templates Use GPT-4o to paraphrase examples into more naturalistic language
Figure 2: Pipeline for generating each reasoning question in InAbHyD.
@@ 3.1 Overview
    Second, we traverse the ontology tree and randomly mark some axioms as hidden. Each hidden axiom is the ground truth.
@@
    Third, for each hidden axiom, we generate observations so that LLMs have enough information to produce ground-truth hypotheses.
@@
    Fourth, we transform the incomplete world model and observations into natural language.
Subtask Reasoning type tested World model Observations Ground-truth hypothesis Proof trees for observations
@@
Table 2: An overview of various subtasks and their details in the simplest case (a height-1 ontology tree with a single hypothesis). Here o denotes an ontology concept (e.g, cat) and o(A) denotes the membership (e.g., cat(Sally) means Sally is a cat). p denotes a property (e.g, cute) and p(A) denotes A has property p (e.g., cute(Sally) means Sally is cute). We only show the proof tree for one observation for simplicity; other observations can be derived similarly. In the proof tree, we show hypotheses in green and axioms in world models in purple.
@@
    Strong accuracy: Hypotheses are strongly correct if they exactly match ground-truth hypotheses (simplest hypothesis).
@@
    Hypothesis quality: A quantitative measurement of the quality of the LLMs’ produced hypotheses based on Occam’s Razor.
@@
3.2 World model generation
In our approach, the world model is an ontology tree in first-order logic. Each node in the ontology tree is a concept (e.g. cat) with properties (e.g. All cats are cute ) and members (e.g. Amy is a cat). Each edge in the ontology tree denotes a subtype relationship (e.g., the ragdoll node being a child of the cat node implies Each ragdoll is a cat).
@@ 3.2 World model generation
We first generate the topology of the ontology tree with the height as a controllable parameter. Each non-leaf node has a set of child nodes. The height of the ontology tree is a variable that characterizes the complexity of the world model. We then make two passes over the tree using level-order traversal. In the first pass, for each node, we allocate a concept name, properties, and members  [^3] The properties can also be negated, e.g., All cats are not slow, which corresponds to the logical form $\forall x(\texttt{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}cat}}(x)\rightarrow\neg\texttt{{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}slow}}(x))$ (negation always takes narrow scope). The number of properties, members, and child nodes follows a zipf distribution with $a=3$.. After the first pass, we have a complete world model with three types of axioms: (1) properties (e.g., All cats are cute), (2) membership relations (e.g., Amy is a cat), (3) subtype relations (e.g., Each ragdoll is a cat). In the second pass, we randomly mark some axioms as hidden. These axioms are the ground-truth hypotheses. After the second pass, we have an incomplete world model.
3.3 Observation generation